OpenStudy (anonymous):

A local Dunkin’ Donuts shop reported that its sales have increased exactly 22% per year for the last 2 years. This year’s sales were $82,500. What were Dunkin' Donuts,sales 2 years ago? (Round each year’s sales to the nearest dollar.)

OpenStudy (anonymous):

$50,193 I think. $82,500 * .78 = $64,350 $64,350 * 0.78 = $50,193

OpenStudy (anonymous):

wait nvm

OpenStudy (anonymous):

55,429.

OpenStudy (anonymous):

82500/1.22= 67623/1.22= 55429